I have a question that states: find the vertical and horizontal asymptotes for the rational function f(x)=(2x^2-5x-3)/(x^2-4)
Answer by venugopalramana(3286)

I have a question that states: find the vertical and horizontal asymptotes for the rational function f(x)= Y =(2x^2-5x-3)/(x^2-4)
FIND LIMIT Y AS X TEND TO INFINITY
LIMIT Y = 2/1=2
HENCE Y=2 IS THE HORIZONTAL ASYMPTOTE.
FIND WHEN Y TENDS TO INFINITY
THIS HAPPENS WHEN D.R=X^2-4=0
X=+2 AND X=-2
THESE ARE THE 2 VERTICAL ASYMPTOTES